A 2.40mH toroidal solenoid has an average radius of 7.00cm and a cross-sectional area of 2.00cm2 A) How many coils does it have? In calculating the flux, assume that B is uniform accross a cross section, neglect the variation of B with distance from the toroidal axis. B) At what rate must the current through it change so that a potential difference of 2.2V is developed across its ends? (answer in A/s)
Answer #1

A. inductance L = u0N^2A/L

N^2 = 0.0024*0.07/(4pi*10^-7 * 2*10^-4)

N^2 = 668790

N = 818 turns


B. induced emf e = NL di/dt

di/dt = 2.2/(0.0024)*818

di/dt = 1.12 A/s
Answer #3

It's 0.0024 instead of 0.024

answered by: Paúl Valdivia